# SmallKV: Small Model Assisted Compensation of KV Cache Compression for Efficient LLM Inference

### v1

@@ Appendix B Extended Discussion of Attention Similarity between Different-scale LLMs
We speculate that there are many factors that cause this similarity between different-scale LLMs within same series: 1) Although they differences in configuration parameters such as the number of layers and hidden dimensions, they share the same model architecture. 2) They are pre-trained on the similar corpus [Yang et al., 2024a]. 3) Some smaller models are distilled from larger models [Team et al., 2024]. The aforementioned factors contribute to different-scale LLMs learning similar semantic space representations at different level in the shallow decode layers. For deeper layers, previous researches [Yang et al., 2024b, He et al., 2024] have indicated that LLMs exhibit more attention redundancy, meaning these layers tend to have simpler attention patterns that are easy to match for the SLM, which can be evidenced by the lighter color in latter part of Figure 8.
@@ Appendix F Case Study
Given the same prompt text in a BBT Test, we visualize the output generated by Qwen2-7B under different KV cache compression methods, including the baseline of full cache, H2[Zhang et al., 2023], PyramidInfer [Yang et al., 2024b], and our SmallKV. The results are shown in Figure 9. With only 10% of the KV cache retained, SmallKV produces outputs that are highly similar to those generated with the full cache, both follow the "Let’s think step by step." instruction and calculates according to the steps provided in the prompt examples. In contrast, H2O and PyramidInfer lose their ability to follow instructions and learn from the given reasoning patterns due to compression. Even worse, they invente new words not present in the original task description during the alphabetical ordering task (e.g., "aspect"). Moreover, H2O produces meaningless repeated content at the end of its answer.
Figure 9: Visualized outputs of one generation example with Qwen2-7B. Results are compared between the baseline model with full cache, H2[Zhang et al., 2023], PyramidInfer [Yang et al., 2024b], and our SmallKV.

### v2

@@ Appendix B Extended Discussion of Attention Similarity between Different-scale LLMs
We speculate that there are many factors that cause this similarity between different-scale LLMs within same series: 1) Although they differences in configuration parameters such as the number of layers and hidden dimensions, they share the same model architecture. 2) They are pre-trained on the similar corpus [42]. 3) Some smaller models are distilled from larger models [36]. The aforementioned factors contribute to different-scale LLMs learning similar semantic space representations at different level in the shallow decode layers. For deeper layers, previous researches [43, 13] have indicated that LLMs exhibit more attention redundancy, meaning these layers tend to have simpler attention patterns that are easy to match for the SLM, which can be evidenced by the lighter color in latter part of Figure 8.
@@
To analyze the attention similarity between different SLMs and LLMs, we first visualize the attention similarity among Qwen2-7B, Qwen2-0.5B, and Llama3-1B on the Wikitext-2 dataset (as shown in Figure 9), where the y-axis represents the cumulative attention score per token across all attention heads. Results show that, due to similar training corpora and linguistic priors, models exhibit certain commonalities in their attention patterns. However, the cosine similarity between Qwen2-7B and Qwen2-0.5B reaches 0.9622, while that between Qwen2-7B and Llama3-1B is only 0.6917. This indicates that attention similarity is significantly higher within the same series models.
Figure 9: Accumulated attention scores of Qwen2-7B, Qwen2-0.5B, and Llama3-1B (Selecting a random segment from the middle of the tokens to avoid attention sink).
We further conduct a comprehensive analysis of attention pattern similarity between SLMs and LLMs across diverse models, scales, benchmarks, and multi-turn dialogue scenarios. Due to attention sinks, we removed the first few tokens when computing similarities to avoid too high similarity (e.g.,  0.99), ensuring meaningful and comparable results. The results are represented in Table  7 and Table  8.
Table 7: Similarity between SLM and LLM across different settings (Part 1)
SLM Qwen2-0.5B Qwen2-0.5B Qwen2-0.5B Qwen3-0.6B Qwen2.5-0.5B Qwen2.5-1.5B
LLM Qwen2-7B Qwen2-7B Qwen2-7B Qwen3-8B Qwen2.5-14B Qwen2.5-14B
Benchmark BBH GSM8K MMLU BBH BBH BBH
Similarity 0.85 0.85 0.98 0.83 0.83 0.83
Table 8: Similarity between SLM and LLM across different settings (Continued)
SLM Qwen2.5-3B Qwen2-7B Qwen2.5-7B LlaMA3-2.1B LLaMA3.2-1B
LLM Qwen2.5-14B Qwen2-72B Qwen2.5-14B Qwen2-7B LLaMA3.1-8B
Benchmark BBH BBH BBH BBH BBH
Similarity 0.85 0.81 0.86 0.56 0.98
Table 9: Attention similarity on multi-turn dialogues.
global turn1 turn2 turn3 turn4 turn5
SmallKV 0.822 0.812 0.799 0.796 0.782
We evaluate the changes in similarity during multi-turn dialogue scenarios, and the results are shown in Table  9. These findings demonstrate the scalability of attention similarities to a considerable extent.
@@ Appendix F Case Study
Given the same prompt text in a BBT Test, we visualize the output generated by Qwen2-7B under different KV cache compression methods, including the baseline of full cache, H2[51], PyramidInfer [43], and our SmallKV. The results are shown in Figure 10. With only 10% of the KV cache retained, SmallKV produces outputs that are highly similar to those generated with the full cache, both follow the "Let’s think step by step." instruction and calculates according to the steps provided in the prompt examples. In contrast, H2O and PyramidInfer lose their ability to follow instructions and learn from the given reasoning patterns due to compression. Even worse, they invente new words not present in the original task description during the alphabetical ordering task (e.g., "aspect"). Moreover, H2O produces meaningless repeated content at the end of its answer.
Figure 10: Visualized outputs of one generation example with Qwen2-7B. Results are compared between the baseline model with full cache, H2[51], PyramidInfer [43], and our SmallKV.